# Robust Majorana bound state in pseudo-spin domain wall of 2-D topological insulator

### 02 Feb 2024 v1

@@ Abstract
We investigate helical edge states (HES) emerging at the composite domain wall of spin and pseudo-spin degrees of freedom in a 2-D bulk governed by the Bernevig-Hughes-Zhang Hamiltonian which underwent quantum spin Hall to anomalous Hall transition. We numerically study the stability of Majorana bound state (MBS) formed due to proximity induced superconductivity in these helical edge states. We establish exceptional robustness of MBS against moderate chemical potential or magnetic disorder owing to the existence of the simultaneous orthogonality between the right and the left moving modes both in spin and pseudo-spin space. Hence our proposal could pave the way to realizing robust Majorana bound state on 2D platforms.
@@
We investigate the transport of electrons across the emergent HES by employing quantum transport in 2D lattice simulations using the Bernevig-Hughes-Zhang (BHZ) model Bernevig et al. (2006). The key results of our numerical simulations can be summarized as follows: (i) Backscattering of electrons in the emergent HES enjoys unprecedented protection against local disorder potential which breaks the spin or the pseudo-spin rotational symmetry individually (ii) The emergent HES when exposed to a proximity-induced superconductivity hosts MBS which show remarkable resilience against chemical potential or time-reversal symmetry breaking magnetic disorder of moderate strength (with respect to the bulk QAH gap).
@@
We impose a domain wall in the exchange field profile given by $g(x)=g_{0}\left[\Theta(L/2-x)-\Theta(x-L/2)\right]$ (with $|g_{0}|>|\mathcal{M}|$), which creates two chiral co-propagating edges with orthogonal spin-polarization at the domain wall. To make these edges counter-propagating like in a HES, we impose an additional domain wall in the parameter $A\left[\Theta(L/2-x)-\Theta(x-L/2)\right]$. Changing the parameter $A$ to $-A$ in the $x-$hopping inverts the propagation direction of electrons along the edge, which means in presence of this domain wall, electrons in the left and right half of the system flow with the same helicity (see App. A for more details). This also gives rise to mutual orthogonality in pseudo-spin states of the two counter-propagating vertical edges at the domain wall which we will discuss in detail in the next section. So, the simultaneous presence of the domain walls in $g(x)$ and $A(x)$ creates orthogonality in spin and pseudo-spin states of the counterpropagating edge states at the domain wall and gives rise to an emergent HES with orthogonality in spin and pseudo-spin space. The tight-binding Hamiltonian in the whole region in the presence of the aforementioned domain walls can be written in a concise form,
@@
Figure 2: Plots depicting (a) probability current for spin-$\uparrow$ electrons injected from the “Lead-1” and (b) probability current for spin-$\downarrow$ electrons injected from the “Lead-2” in absence of any disorder potential at the junction region of “QAH-1” and “QAH-2”. The arrows in the plots signify the propagation direction of injected electrons which shows that the spin-$\uparrow$ and spin-$\downarrow$ electrons injected from “Lead-1” and “Lead-2” are counter-propagating at the domain wall. The horizontal and vertical axes of the rectangular region represent the actual physical dimension of the system in nm.
Figure 3: Plots depicting (a) spin ($z$-component) (b) pseudospin ($x$-component) (c) pseudospin ($y$-component) densities of “QAH-1” for injected spin-$\uparrow$ electrons from the “Lead-1” and (d) spin ($z$-component) (e) pseudospin ($x$-component) (f) pseudospin ($y$-component) densities of “QAH-2” for injected spin-$\downarrow$ electrons from the “Lead-2” in absence of any disorder potential at the junction region of “QAH-1” and “QAH-2”. The horizontal and vertical axes of the rectangular region represent the actual physical dimension of the system in nm.
@@
In Fig. 3-(b), (c), (e), and (f), pseudo-spin polarization of the edge states in QAH-1 and QAH-2 are shown. The pseudo-spin polarizations of the top edges in QAH-1 and QAH-2 are oriented along the $-\hat{x}$ direction, while for the bottom edges, the pseudospin polarization is oriented along $\hat{x}$. The pseudo-spins of the vertical edges in the two QAH regions are $-\hat{y}$ and $\hat{y}$ polarized, respectively. Due to the inherent block-diagonal structure of a standard BHZ Hamiltonian in momentum space, the edge state vectors of the QAH-1 and QAH-2 edges can be elegantly expressed as the direct product of spin and pseudo-spin states, respectivelyRoy et al. (2016). We denote the vertical edge states of QAH-$i$ at the domain wall as $|V_{i}\rangle$ (for $i=1,2$) and building upon the preceding analysis, the spinors can be expressed as (see App. A),
@@ 4 $5(\sigma_{x}+\tilde{\sigma}_{x})$ 0 1 0
To investigate the transport characteristics across the emergent HES, we carry out transport simulation such that spin-$\uparrow$ electrons are injected from “Lead-1” and we observe complete electron-to-electron reflection; no electron transmission from “Lead-1” to “Lead-2” lead as expected due to the orthogonality present in both the spin and pseudo-spin states of two edge channels $|V_{1}\rangle$ and $|V_{2}\rangle$ of the emergent HES. Also, it is evident that at the domain wall, applying any kind of edge local potential term of type $\sigma_{x}\tilde{\sigma}_{x}$, $\sigma_{x}\tilde{\sigma}_{z}$, $\sigma_{y}\tilde{\sigma}_{x}$, $\sigma_{y}\tilde{\sigma}_{z}$ that simultaneously couples both the spin and pseudo-spin states of $|V_{1}\rangle$ and $|V_{2}\rangle$ should results in the opening of a gap in the emergent HES by hybridizing these states. Consequently, spin-$\uparrow$ electrons injected from “Lead-1” should transmit perfectly into “Lead-2” as spin-$\downarrow$ electrons. On the other hand, any potential which couples to only spin or pseudo-spin state should be of no consequence. Transmission of electrons in the presence and absence of various edge potentials in the emergent HES obtained from our transport simulations is presented in a tabular form in Table. I, which support the above discussed. In the above discussion, the applied potentials on the edge follow a spatial profile $\Theta(L/2-x)\Theta(x-(L/2-1))$, which means it is present only at the domain wall where $L$ is defined in terms of lattice spacing is given in the previous section. In what follows we study the transport across the emergent HES in the presence of a local superconducting pairing term.
$+\Delta\Theta(i_{x}-(i_{x_{c}}-5))\Theta(i_{x_{c}}-i_{x})\tau_{x},$
$H_{i,i+a_{x}}$ $=\frac{D+B\tilde{\sigma}_{z}}{a^{2}}\tau_{z}$
@@
Table 2: Our investigation of transport across the superconductor proximitized emergent HES, both in the presence and absence of various edge local potentials at zero energy ($E=0$ meV), is shown. We apply different kinds of edge local potential terms denoted as $M_{x}$ on the emergent HES and examine their effects on the transport of electrons. Throughout our discussions, we represent particle-to-particle transmission from “Lead-1” to “Lead-1” as reflection and “Lead-1” to “Lead-2” as transmission. We express the electron-to-electron reflection, electron-to-hole reflection, electron-to-electron transmission, and electron-to-hole transmission probabilities as $R_{ee}$, $R_{he}$, $T_{ee}$, and $T_{he}$, respectively. Assuming electron injection from “Lead-1”, the Andreev conductance in “Lead-1” is expressed as $G=(e^{2}/h)(1-T_{ee}+T_{he})$. All relevant energy units are represented in meV. At critical point $M_{x}=\Delta$, the charge current becomes zero, although the spin-current remains non-zero.
The Nambu basis consider here is given by, $\Psi=(\psi_{1\uparrow},\psi^{\dagger}_{1\downarrow},\psi_{2\uparrow},\psi^{\dagger}_{2\downarrow},\psi_{1\downarrow},-\psi^{\dagger}_{1\uparrow},\psi_{2\downarrow},-\psi^{\dagger}_{2\uparrow})$ where $1,2$ represent the orbitals, and $\uparrow$,$\downarrow$ , ↓ represent the spin indices along the $S_{z}$ direction. The plots depicting the transport of electrons across the superconductivity proximitized HES as a function of the ratio of the superconducting gap ($\Delta$) and the edge potential($M_{x}$) are shown in Fig. 4, where $R_{ee}$, $R_{he}$. $T_{ee}$ and $T_{he}$ denote the electron-to-electron reflection, electron-to-hole reflection, electron-to-electron transmission, and electron-to-hole transmission probabilities respectively between the left and right lead. We define the local conductance in “Lead-1” as $G_{11}=(N_{e}+R_{he}-R_{ee})e^{2}/h$, where $N_{e}$ is the number of channels for electron injection in “Lead-1” and the non-local conductance in “Lead-2” as $G_{21}=(T_{ee}-T_{he})$ Fuchs et al. (2021); Zhang and Trauzettel (2019); Haugen et al. (2010). In the presence of the superconducting pairing term $\Delta$, the spin-$\uparrow$ electron injected from “QAH-1” pairs up with another spin-$\downarrow$ electron from “QAH-2” to form a cooper pair and the spin-$\downarrow$ electron leaves behind a spin-$\downarrow$ hole which is received at the “Lead-2”. Applying the edge local potential term $M_{x}\sigma_{x}\tilde{\sigma}_{x}$ in the region where local superconductivity has been induced, causes the emergent HES to coherently backscatter, thereby trying to open an insulating gap in the edge state. This gap competes with the superconducting gap, hence leading to suppression in $T_{he}$. When the edge local potential term becomes equal to the superconducting gap, reaching a gap-closing point (critical point), every type of possible transmission becomes equally probable, viz., $R_{ee}=R_{he}=T_{ee}=T_{he}=0.25$. A similar kind of transport where all transmission and reflection probabilities become equal to $1/4$ was also observed in a single channel $\mathcal{T}$-stub geometry strongly coupled to a superconductor at resonant energies of incidenceDas et al. (2009). At this critical point, the non-local electrical conductance $G_{21}=(e^{2}/h)(T_{ee}-T_{he})$ Fuchs et al. (2021); Zhang and Trauzettel (2019); Haugen et al. (2010), becomes zero which means zero average charge current is detected at “Lead-2”, while the average spin current in the lead should remain finite. As we cross the critical point, i.e., when the insulating gap exceeds the superconducting gap, we observe a rise in electron-to-electron transmission between the two leads. The Andreev conductance in “Lead-1” (assuming electron injection from “Lead-1”) is expressed as $G=(1-T_{ee}+T_{he})e^{2}/h$ Fuchs et al. (2021); Adak et al. (2022); Falci et al. (2001) and the $2e^{2}/h$ Andreev conductance peak (when $T_{he}=1,T_{ee}=0$) at zero energy signifies the presence of MBS.
Figure 4: Plots depicting (a) $R_{ee}$ vs $M_{x}/\Delta$ (b) $R_{he}$ vs $M_{x}/\Delta$ (c) $T_{ee}$ vs $M_{x}/\Delta$ (d) $T_{he}$ vs $M_{x}/\Delta$ behaviour for electrons with incident energies $E$ (expressed in meV), keeping $\Delta$ fixed to $5$ meV. The potential ($M_{x}$) applied on the HES is of type $\sigma_{x}\tilde{\sigma}_{x}$. The faint black line represents the $0.25$ probability value. The non-local electrical conductance as measured in the “Lead-2” is given by $G_{21}=(e^{2}/h)(T_{ee}-T_{he})$ Fuchs et al. (2021); Zhang and Trauzettel (2019); Haugen et al. (2010). The plots show that at the critical point ($M_{x}=\Delta$, where the gap in the HES spectrum closes), the net electrical conductance measured in the “Lead-2” is zero.
These phenomena resemble the induction of superconductivity by proximity to a conventional helical edge, causing the edge spectrum to open a gap and host MBS at the end of proximitized region Alicea (2012). However, the advantage of this emergent HES over conventional HES lies in the fact that the MBS hosted by the emergent HES remains remarkably robust against moderate chemical potential and time-reversal symmetry-breaking magnetic disorder, which we discuss in detail in the next section. We present the study of transport across the emergent HES in the presence of local superconductivity and edge local potentials at the emergent HES in a tabular form in Table. II.
Taken together, these findings illuminate the potential for realizing resilient MBS at the HES emerging at the spin and pseudo-spin domain wall within the 2D topological insulator, owing to the presence of the simultaneous spin and pseudo-spin orthogonality.
Conclusion : In conclusion, our study offers compelling numerical evidence through a lattice model simulation of a 2D topological insulator, that the presence of the proposed domain wall leads to the emergence of a HES protected by spin and pseudo-spin orthogonality. This emergent HES in the presence of proximity induced superconductivity hosts MBS that exhibit remarkable resilience against chemical potential and magnetic disorders present in the system. As a result, this model presents a promising platform to realize disorder robust effective HES and Majorana zero modes.
Appendix A Spin and pseudo-spin polarizations in presence of the domain walls
The structure of $h_{u}$ and $h_{d}$ are identical, and that is why, the spin-$\uparrow$ electrons in “QAH-1” and spin-$\downarrow$ electrons in “QAH-2” flow with a similar helicity. So one can conveniently write down the vertical edge state vector of the right half region (“QAH-2”) as,
As a result, at the domain wall, we have two oppositely spin-polarized counter-propagating edge states with orthogonal pseudo-spin polarization (see Fig. 3) which gives rise to an emergent HES.

### 07 Mar 2024 v2

@@ Abstract
We investigate helical edge states (HES) emerging at the composite domain wall of spin and pseudo-spin degrees of freedom in a 2-D bulk governed by the Bernevig-Hughes-Zhang Hamiltonian which underwent quantum spin Hall to anomalous Hall transition. We numerically study the stability of Majorana bound state (MBS) formed due to proximity induced superconductivity in these helical edge states. We establish exceptional robustness of MBS against moderate chemical potential or magnetic disorder owing to the existence of the simultaneous orthogonality between the right and the left moving modes both in spin and pseudo-spin space. Hence our proposal could pave the way to realizing robust Majorana bound state on 2-D platforms.
@@
We investigate the transport of electrons across the emergent HES by employing quantum transport in 2-D lattice simulations using the Bernevig-Hughes-Zhang (BHZ) model Bernevig et al. (2006). The key results of our numerical simulations can be summarized as follows: (i) Backscattering of electrons in the emergent HES enjoys unprecedented protection against local disorder potential which breaks the spin or the pseudo-spin rotational symmetry individually (ii) The emergent HES when exposed to a proximity-induced superconductivity hosts MBS which show remarkable resilience against chemical potential or time-reversal symmetry breaking magnetic disorder of moderate strength (with respect to the bulk QAH gap).
@@
We impose a domain wall in the exchange field profile given by $g(x)=g_{0}\left[\Theta(L/2-x)-\Theta(x-L/2)\right]$ (with $|g_{0}|>|\mathcal{M}|$), which creates two chiral co-propagating edges with orthogonal spin-polarization at the domain wall. To make these edges counter-propagating like in a HES, we impose an additional domain wall in the parameter $A\left[\Theta(L/2-x)-\Theta(x-L/2)\right]$ (see Eq. 3). Changing the parameter $A$ to $-A$ in the $x-$hopping inverts the propagation direction of electrons along the edge, which means in presence of this domain wall, electrons in the left and right half of the system flow with the same helicity (see App. A for more details). This also gives rise to mutual orthogonality in pseudo-spin states of the two counter-propagating vertical edges at the domain wall which we will discuss in detail in the next section. So, the simultaneous presence of the domain walls in $g(x)$ and $A(x)$ creates orthogonality in spin and pseudo-spin states of the counterpropagating edge states at the domain wall and gives rise to an emergent HES with orthogonality in spin and pseudo-spin space. The tight-binding Hamiltonian in the whole region in the presence of the aforementioned domain walls can be written in a concise form,
@@
Figure 2: Plots depicting (a) probability current for spin-$\uparrow$ electrons injected from the “Lead-1” and (b) probability current for spin-$\downarrow$ electrons injected from the “Lead-2” in absence of any disorder potential at the junction region of “QAH-1” and “QAH-2”. The arrows in the plots denote the propagation direction of injected electrons which shows that the spin-$\uparrow$ and spin-$\downarrow$ electrons injected from “Lead-1” and “Lead-2” are counter-propagating at the domain wall (junction region of “QAH-1” and “QAH-2”). The horizontal and vertical axes of the rectangular region represent the actual physical dimension of the system in nm.
Figure 3: Plots depicting (a) spin ($z$-component) (b) pseudospin ($x$-component) (c) pseudospin ($y$-component) densities of “QAH-1” for injected spin-$\uparrow$ electrons from the “Lead-1” and (d) spin ($z$-component) (e) pseudospin ($x$-component) (f) pseudospin ($y$-component) densities of “QAH-2” for injected spin-$\downarrow$ electrons from the “Lead-2” in absence of any disorder potential at the domain wall (junction region of “QAH-1” and “QAH-2”). The horizontal and vertical axes of the rectangular region represent the actual physical dimension of the system in nm.
@@
In Fig. 3-(b), (c), (e), and (f), pseudo-spin polarization of the edge states in QAH-1 and QAH-2 are shown. The pseudo-spin polarizations of the top edges in QAH-1 and QAH-2 are oriented along the $-\hat{x}$ direction, while for the bottom edges, the pseudospin polarization is oriented along $\hat{x}$. The pseudo-spins of the vertical edges in the two QAH regions are $-\hat{y}$ and $\hat{y}$ polarized, respectively. Due to the inherent block-diagonal structure of a standard BHZ Hamiltonian in momentum space, the edge state vectors of the QAH-1 and QAH-2 edges can be elegantly expressed as the direct product of spin and pseudo-spin states, respectivelyRoy et al. (2016). We denote the vertical edge states of QAH-$i$ at the domain wall as $|V_{i}\rangle$ (for $i=1,2$) and building upon the preceding analysis, the spinors can be expressed as (see Supplemental Material Section A),
@@
To investigate the transport characteristics across the emergent HES, we carry out transport simulation such that spin-$\uparrow$ electrons are injected from “Lead-1” and we observe complete electron-to-electron reflection; no electron transmission from “Lead-1” to “Lead-2” lead as expected due to the orthogonality present in both the spin and pseudo-spin states of two edge channels $|V_{1}\rangle$ and $|V_{2}\rangle$ of the emergent HES. Also, it is evident that at the domain wall, applying any kind of edge local potential term of type $\sigma_{x}\tilde{\sigma}_{x}$, $\sigma_{x}\tilde{\sigma}_{z}$, $\sigma_{y}\tilde{\sigma}_{x}$, $\sigma_{y}\tilde{\sigma}_{z}$ that simultaneously couples both the spin and pseudo-spin states of $|V_{1}\rangle$ and $|V_{2}\rangle$ should results in the opening of a gap in the emergent HES by hybridizing these states. Consequently, spin-$\uparrow$ electrons injected from “Lead-1” should transmit perfectly into “Lead-2” as spin-$\downarrow$ electrons. On the other hand, any potential which couples to only spin or pseudo-spin state should be of no consequence. Transmission of electrons in the presence and absence of various edge potentials in the emergent HES obtained from our transport simulations is presented in a tabular form in Table. I in Supplemental Material Section B, which supports the above discussed. In the above discussion, the applied potentials on the edge follow a spatial profile $\Theta(L/2-x)\Theta(x-(L/2-1))$, which means it is present only at the domain wall where $L$ is defined in terms of lattice spacing is given in the previous section. In what follows we study the transport across the emergent HES in the presence of a local superconducting pairing term.
$+\Delta\Theta(i_{x}-(i_{x_{c}}-5))\Theta(i_{x_{c}}-i_{x})\tau_{x},$
$H_{i,i+a_{x}}$ $=\frac{D+B\tilde{\sigma}_{z}}{a^{2}}\tau_{z}$
Figure 4: Plots depicting (a) $R_{ee}$ vs $M_{x}/\Delta$ (b) $R_{he}$ vs $M_{x}/\Delta$ (c) $T_{ee}$ vs $M_{x}/\Delta$ (d) $T_{he}$ vs $M_{x}/\Delta$ behaviour for electrons with incident energies $E$ (expressed in meV), keeping $\Delta$ fixed to $5$ meV. The potential ($M_{x}$) applied on the HES is of type $\sigma_{x}\tilde{\sigma}_{x}$. The faint black line represents the $0.25$ probability value. The non-local electrical conductance as measured in the “Lead-2” is given by $G_{21}=(e^{2}/h)(T_{ee}-T_{he})$ Fuchs et al. (2021); Zhang and Trauzettel (2019); Haugen et al. (2010). The plots show that at the critical point ($M_{x}=\Delta$, where the gap in the HES spectrum closes), the net electrical conductance measured in the “Lead-2” is zero. As a consequence, zero average charge current is received at “Lead-2”, however, the average spin current in “Lead-2” remains non-zero.
The Nambu basis consider here is given by, $\Psi=(\psi_{1\uparrow},\psi^{\dagger}_{1\downarrow},\psi_{2\uparrow},\psi^{\dagger}_{2\downarrow},\psi_{1\downarrow},-\psi^{\dagger}_{1\uparrow},\psi_{2\downarrow},-\psi^{\dagger}_{2\uparrow})$ where $1,2$ represent the orbitals, and $\uparrow$,$\downarrow$ , ↓ represent the spin indices along the $S_{z}$ direction. The plots depicting the transport of electrons across the superconductivity proximitized HES as a function of the ratio of the superconducting gap ($\Delta$) and the edge potential ($M_{x}$) are shown in Fig. 4, where $R_{ee}$, $R_{he}$. $T_{ee}$ and $T_{he}$ denote the electron-to-electron reflection, electron-to-hole reflection, electron-to-electron transmission, and electron-to-hole transmission probabilities respectively between the left and right lead. We define the local conductance in “Lead-1” as $G_{11}=(N_{e}+R_{he}-R_{ee})e^{2}/h$, where $N_{e}$ is the number of channels for electron injection in “Lead-1” and the non-local conductance in “Lead-2” as $G_{21}=(T_{ee}-T_{he})$ Fuchs et al. (2021); Zhang and Trauzettel (2019); Haugen et al. (2010). In the presence of the superconducting pairing term $\Delta$, the spin-$\uparrow$ electron injected from “QAH-1” pairs up with another spin-$\downarrow$ electron from “QAH-2” to form a cooper pair and the spin-$\downarrow$ electron leaves behind a spin-$\downarrow$ hole which is received at the “Lead-2”. Applying the edge local potential term $M_{x}\sigma_{x}\tilde{\sigma}_{x}$ in the region where local superconductivity has been induced, causes the emergent HES to coherently backscatter, thereby trying to open an insulating gap in the edge state. This gap competes with the superconducting gap, hence leading to suppression in $T_{he}$. When the edge local potential term becomes equal to the superconducting gap, reaching a gap-closing point (critical point), every type of possible transmission becomes equally probable, viz., $R_{ee}=R_{he}=T_{ee}=T_{he}=0.25$. A similar kind of transport where all transmission and reflection probabilities become equal to $1/4$ was also observed in a single channel $\mathcal{T}$-stub geometry strongly coupled to a superconductor at resonant energies of incidenceDas et al. (2009). At this critical point, the non-local electrical conductance $G_{21}=(e^{2}/h)(T_{ee}-T_{he})$ Fuchs et al. (2021); Zhang and Trauzettel (2019); Haugen et al. (2010), becomes zero which means zero average charge current is detected at “Lead-2”, while the average spin current in the lead should remain non-zero. As we cross the critical point, i.e., when the insulating gap exceeds the superconducting gap, we observe a rise in electron-to-electron transmission between the two leads. The Andreev conductance in “Lead-1” (assuming electron injection from “Lead-1”) is expressed as $G=(1-T_{ee}+T_{he})e^{2}/h$ Fuchs et al. (2021); Adak et al. (2022); Falci et al. (2001) and the $2e^{2}/h$ Andreev conductance peak (when $T_{he}=1,T_{ee}=0$) at zero energy signifies the presence of MBS.
These phenomena resemble the induction of superconductivity by proximity to a conventional helical edge, causing the edge spectrum to open a gap and host MBS at the end of proximitized region Alicea (2012). However, the advantage of this emergent HES over conventional HES lies in the fact that the MBS hosted by the emergent HES remains remarkably robust against moderate chemical potential and time-reversal symmetry-breaking magnetic disorder, which we discuss in detail in the next section. We present the study of transport across the emergent HES in the presence of local superconductivity and edge local potentials at the emergent HES in a tabular form in Table. II in Supplemental Material Section B.
Conclusion : In conclusion, our study offers compelling numerical evidence through a lattice model simulation of a 2-D topological insulator, that the presence of the proposed domain wall leads to the emergence of an HES protected by spin and pseudo-spin orthogonality. This emergent HES in the presence of proximity induced superconductivity hosts MBS that exhibit remarkable resilience against chemical potential and magnetic disorders present in the system. As a result, this model presents a promising platform to realize disorder robust effective HES and Majorana zero modes.
Supplemental Material A Spin and pseudo-spin polarizations in presence of the domain walls
The structure of $h_{u}$ and $h_{d}$ are identical, and that is why, the spin-$\uparrow$ electrons in “QAH-1” and spin-$\downarrow$ electrons in “QAH-2” flow with a similar helicity (in a clockwise fashion as shown in Fig. 1 of the main text). So one can conveniently write down the vertical edge state vector of the right half region (“QAH-2”) at the domain wall as,
where $f(-x)$ is a spatially decaying part which becomes zero at $x\rightarrow+\infty$.
As a result, at the domain wall, we have two oppositely spin-polarized counter-propagating edge states with orthogonal pseudo-spin polarization (see Fig. 2 and Fig. 3 of the main text) which gives rise to an emergent HES.
Supplemental Material B Transport of electrons at zero energy across the emergent HES
Below we present our study of transport of electrons at zero energy across the emergent HES both in presence and absence of proximitized local superconductivity ($\Delta$) and edge local potentials ($M_{x}$) at the emergent HES.
@@ 4 $5(\sigma_{x}+\tilde{\sigma}_{x})$ 0 1 0
Table 2: Our investigation of transport across the superconductor proximitized emergent HES, both in the presence and absence of various edge local potentials at zero energy ($E=0$ meV), is shown. We apply different kinds of edge local potential terms denoted as $M_{x}$ on the emergent HES and examine their effects on the transport of electrons. Throughout our discussions, we represent particle-to-particle transmission from “Lead-1” to “Lead-1” as reflection and “Lead-1” to “Lead-2” as transmission. We express the electron-to-electron reflection, electron-to-hole reflection, electron-to-electron transmission, and electron-to-hole transmission probabilities as $R_{ee}$, $R_{he}$, $T_{ee}$, and $T_{he}$, respectively. Assuming electron injection from “Lead-1”, the Andreev conductance in “Lead-1” is expressed as $G=(e^{2}/h)(1-T_{ee}+T_{he})$. All relevant energy units are represented in meV. At critical point $M_{x}=\Delta$, the average charge current detected in “Lead-2” becomes zero, although the spin-current remains non-zero.